Wave Crest Cosmetics produces cosmetic products for 11 major buyers. One customer accounts for ½ of the sales, the other 10 customers account for the remaining sales. The 10 smaller customers place orders of about equal size. The large customer purchases 1,000,000 units per year by placing 3 orders per year, and the 10 smaller customers order 1,000,000 units in total by placing 400 orders per year. Order filling costs total $685,100 per year. How much would be assigned to the large customer as ordering costs using activity-based customer costs? a. $5,100 b. $342,550 c. $4,700 d. $680,000
Answer #1

Order filling cost per order

= 685,100/(3+400) = 1700 per order

Order filling cost allocated to large customers

= 1700 * 3

= 5100